OpenStudy (anonymous):

solve using the quadratic formula 3x^2-x+6=0

OpenStudy (anonymous):

\[\frac{ 1 \pm \sqrt{72}i }{ 6 }\] Is what I got.

OpenStudy (anonymous):

You simply plug in the numbers, a = 3, b = -1, c =6

OpenStudy (anonymous):

i know but i keep on getting 73 instead of 72

OpenStudy (anonymous):

When you get \[\sqrt{1-72} = \sqrt{71}i\]

OpenStudy (anonymous):

\[\sqrt{(-1)^2 - 4*3*6}\]

OpenStudy (anonymous):

ohh kay